OpenStudy (anonymous):

solve the system: y= -1/3 + 2 and x + 3y = 3

OpenStudy (anonymous):

-1/3x*

OpenStudy (anonymous):

x+3y=3 substrct x from both sides what would you get?

OpenStudy (anonymous):

3y=-x+3?

OpenStudy (anonymous):

id set them both equal to y then solve like that

OpenStudy (anonymous):

ar you can sub y=-1/3x+2 in to other equation x+3(-1/3 x+x)=3 solve for x

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

since y equals -1/3x+2 we can substitute that into the second equation like this:\[x+3(-1/3 x+2)=3\]and then solve for x:\[x-3x+6=3\]\[-2x+6=3\]\[-2x=-3\]\[x=3/2\] now substitute this x into the first equation:\[y=-1/3(3/2)+2\]\[y=-1/2+2\]\[y=-5/2\] \[{x=3/2~~~~~~~~y=-5/2}\]

OpenStudy (anonymous):

okay I see the first part. but the problem with -5/2, is that thats not one of the answer options. but 3/2, 0 is

OpenStudy (anonymous):

oh sorry 1 second

OpenStudy (anonymous):

\[x+3(−1/3x+2)=3\] and then solve for x: \[x−x+6=3\]\[0x+6=3\]\[0x=−3\]\[x=0\] now substitute this x into the first equation: \[y=−1/3(0)+2\]\[y=0+2\]\[y=2\] \[x=0~~~y=2\]

OpenStudy (anonymous):

is that one of the choices? sorry i did that wrong but i fixed it...helpful? :)

OpenStudy (anonymous):

(0, 1) (1, 0) (3, 1/3) (3/2, 0) these are the options.

OpenStudy (anonymous):

im so confused .-.

OpenStudy (anonymous):

its A

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

y has to equal 1 when x is 0

jimthompson5910 (jim_thompson5910):

yummydum, if 0x=−3, then 0 = -3 which is a contradiction So there are no solutions.

OpenStudy (anonymous):

besides the mistake then you .. wait if you have choses then the order pair are (x,y) sub into the equations example take the first one 0,1 y= -1/3 + 2 and x + 3y = 3 1=-1/3(0)+2 and 0+3(1)=3 1=2 not a solutation and 3=3 solves right equations so do the next set of pairs and tell me what you find out

OpenStudy (anonymous):

damn im confused...

OpenStudy (anonymous):

is that all the possiablities? there is no solutions be these are parallal lines

OpenStudy (anonymous):

sorry your confused but it says solve so look at what we have sub the points in the equations and see if it equals or solve for x or y and then find what you did nt solve for... do you understand this

OpenStudy (anonymous):

there is a no solution

jimthompson5910 (jim_thompson5910):

i would be confused too...the answer choices are incorrect because the true answer is "no solution" or "there is/are no solution(s)"

jimthompson5910 (jim_thompson5910):

so there must be a typo somewhere

OpenStudy (anonymous):

there is your answer but do you understand why?

OpenStudy (anonymous):

i think i got this jim_thompson5910

OpenStudy (anonymous):

yeah i do thanks :)

OpenStudy (anonymous):

great you also can see let equations equal -1/3 x+2=-1/3 x+1and solve for x then you get 2=1 so no solution. :-)

OpenStudy (phi):

another way to look at \[y= -\frac{1}{3}x + 2 \text { and } x + 3y = 3\] the 2nd equation, after rearranging into y= mx+b form is \[ y= -\frac{1}{3}x+3\] You have two lines that are parallel, and never meet. There is no (x,y) pair that is on both lines (as would be the case if they intersected)

OpenStudy (anonymous):

already said this phi